### v1

# A new expansion of the Riemann zeta function

Abstract
After a brief introduction to Ramanujan’s method of summation, we give an expansion of the Riemann Zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$ where the functions $P_{m}$ are polynomials with their roots on the line $\{\Re(s)=1/2\}$, the coefficients $x_{m}$ being finite linear combinations of the Euler constant $\gamma$ and the values $\zeta(2),\zeta(3),\dots,\zeta(m+1).$
1 Introduction
There are many polynomial expansions of the Riemann zeta function. Some of them (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}c_{m}\binom{s-1}{m}$,
with
$c_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-1)^{k}\big(\zeta(k+1)-\frac{1}{k}\big).$
Others are linked to the well known expansion
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}\gamma_{m}\frac{(s-1)^{m}}{m!}$
where $\gamma_{m}$ are the Stieltjes constants. This gives (cf.[E-G])
$\zeta(s)-\frac{s}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}l_{m}(\frac{s-1}{s})^{m},$
with
$l_{m}=\sum_{k=1}^{m}\binom{m-1}{k-1}(-1)^{m-k}\frac{\gamma_{k}}{k!}.$
and for $s=\frac{1}{2}+it$ $(t\in\mathbb{R}$) one obtain an expansion of $\zeta(\frac{1}{2}+it)$ on the orthonormal basis $(e_{m})$ of $L^{2}(\mathbb{R},\frac{dt}{2\pi(t^{2}+1/4)})$ given by
$e_{m}(t)=\Big(\frac{1-2it}{1+2it}\Big)^{m}.$
There is also (cf. [K]) an expansion of the function
$\Xi(t)=\frac{s(s-1)}{2}\pi^{-s/2}\Gamma(s/2)\zeta(s),\ \ \ s=\frac{1}{2}+it,\ (t\in\mathbb{R}),$
which takes the form
$\Xi(t)=e^{\pi t/4}\sum_{m=0}^{+\infty}b_{m}p_{m}(t),$
where $p_{m}(t)$ are polynomials related to the Mellin transform of odd Hermite functions. These polynomials are orthogonal in $(L^{2}(\mathbb{R},|\Gamma(3/4+it/2)|^{2}dt)$, but it seems that the coefficients $b_{m}$ does not have a simple form.
The polynomial expansion we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ are the Laguerre polynomials defined by
$L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\frac{x^{k}}{k!}.$
The generating function of the Laguerre polynomials (cf. [Leb]) gives for $x\in[0,+\infty[$
$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\psi_{m}(z).$
where $(\psi_{m})_{m\geq 0}$ the sequence of the functions defined in the half-plane $\{Re(z)>0\}$ by
$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}.$
By Mellin transform of the function $x\mapsto e^{-xz}$, we obtain for $\Re(z)>0$ and $\Re(s)>0$
$\Gamma(s)z^{-s}=\sum_{m=0}^{+\infty}\Big(2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx\Big)\psi_{m}(z),$
Since we can write
$2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx={\Gamma(s)}P_{m}(s),$
where $P_{m}$ is the polynomial
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{\Gamma(s+k)}{\Gamma(s)}\frac{1}{k!},$
then we have, for $\Re(s)>0$, the expansion
$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)P_{m}(s).$
Take $z=1,2,3,\dots$, then we get an infinite sequence of relations that it’s tempting to sum up, to obtain
$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\psi_{m}(n)\Big)P_{m}(s).$
In section 3, we prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
In section 2 we give a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1},$
then, after an explicit evaluation of $z_{n}$, we obtain, for $0<\Re(s)<1$, the following expansion
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$
with
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$
and $r_{m}$ are rational numbers explicitely given. For $s=\frac{1}{2}+it$ with $t\in\mathbb{R}$, the preceding relation gives the expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
where the polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
form complete orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. Thus, for every $m\geq 1$, the roots of $Q_{m}$ are real, and this implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
2 Ramanujan’s summation
2.1 Ramanujan’s constant of a series
Let a function $f\in C^{1}(]0,+\infty[)$, the first order Euler-MacLaurin formula says that, for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}\big(x-[x]-\frac{1}{2}\big)f^{\prime}(x)\ dx$}, (1)
where $[x]$ denote the integer part of $x$. For $x\geq 1$, let
$b(x)=x-[x]-\frac{1}{2}$
and let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+C(f)+\frac{1}{2}f(n)-\int_{n}^{+\infty}b(x)f^{\prime}(x)\ dx$}, (2)
 $\text{ with }\ \ C(f)=\frac{1}{2}f(1)+\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx.$
The constant $C(f)$ is called Ramanujan’s constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
Clearly $C(f)$ depends linearly of $f$. And if the series $\sum_{n\geq 1}f(n)$ is convergent, then by (2) we have
$C(f)=\sum_{k=1}^{+\infty}f(k)-\int_{1}^{+\infty}f(x)\ dx.$
We observe that $C(f)$ is also well defined for some divergent series. We have immediately $C(1)=1/2.$ And applying (2) to the function $f:x\mapsto 1/x$, we have for every integer $n\geq 1$
$\sum_{k=1}^{n}\frac{1}{k}=\log n+C(f)+\frac{1}{2n}+\int_{n}^{+\infty}b(x)\frac{1}{x^{2}}\ dx,$
which gives that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
$C(f)=\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
2.2 Plana’s formula
The following summation formula was discovered by Plana in 1820 (and also given by Abel in 1823) (cf. [Lin]).
Theorem 1.
Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $C_{c,d}>0$ with
$\left|f(x+it)\right|\leq C_{c,d}\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and all }t\in\mathbb{R}.$
Then we have Plana’s formula : for any integer $n\geq 1$,
$\sum_{k=1}^{n}f(k)$ $=\frac{f(1)+f(n)}{2}+\!\int_{1}^{n}f(u)\ du$
$+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt-i\!\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (3)
If we now compare Euler-MacLaurin formula (1) and Plana’s (3), then we get, for $n\geq 1$,
$\int_{1}^{n}b(x){f^{\prime}(x)}\ dx=i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$
If the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\ dt=0,$
then we have
$\int_{1}^{+\infty}b(t){f^{\prime}(t)}\ dt=i\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt.$
Thus, in this case the Ramanujan’s constant $C(f)$, previously defined by (2), is also
$C(f)=\frac{1}{2}f(1)+i\int_{0}^{+\infty}\frac{{f(1+it)-f(1-it)}}{e^{2\pi t}-1}\,\ dt.$ (4)
Take now this last formula as the definition of Ramanujan’s constant $C(f)$ for a function $f$ verifying the hypotheses of theorem 1.
Remark 1.
Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$. Then Ramanujan’s constants $C(f)$ and $C(g)$, defined by (4), need not necessarily be the same. For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
$C(g)=i\int_{0}^{+\infty}\frac{\sin\big(\pi(1+it)\big)-\sin\big(\pi(1-it)\big)}{e^{2\pi t}-1}\,\ dt=\int_{0}^{+\infty}\frac{e^{\pi t}-e^{-\pi t}}{e^{2\pi t}-1}\,\ dt\;\;=\;\int_{0}^{+\infty}e^{-\pi t}\ dt=\frac{1}{\pi}.$
To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that Ramanujan’s constant, defined by (4), depends only on the values $f(n)$ for all integers $n\geq 1$}.
2.3 The space $\mathcal{E}^{\pi}$
Definition 1.
Let ${\alpha}>0$ and ${\mathcal{E}}^{\alpha}$ the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\alpha$ and $\emph{b}>0$ with :
for every ${\varepsilon}>0$, there is a constant $C_{\varepsilon}$ such that
$\left|f(z)\right|\;\;\leq\;\;C_{\varepsilon}\ e^{b\,\Re(z)}\ e^{a\,|\Im(z)|},\ \text{ if }\ \Re(z)>{\varepsilon}.$ (5)
Example 1.
Let $f_{s}\;:\;z\mapsto e^{sz}$, with $s=\sigma+i\tau$, $\sigma\in\mathbb{R}$ and $\tau\in\mathbb{R}$ . For $\Re(z)>0$, we have
$|f_{s}(z)|\;\;=\,\;e^{\sigma\Re(z)-\tau\Im(z)}.$
Thus $f_{s}\in\mathcal{E}^{\alpha}$ if and only if $\alpha<\Im(s)<\alpha$.
Properties
a) If $f\in\mathcal{E}^{\alpha}$ and $g\in\mathcal{E}^{\beta}$, alors $fg\in E^{\alpha+\beta}$.
b) If $f\in\mathcal{E}^{\alpha}$ and if $F(z)=\int_{1}^{z}f(u)du$ then $F\in\mathcal{E}^{\alpha}$.
An important theorem for the rest of this section is the following.
Theorem 2.
Carlson’s theorem (cf. [H2])
If $f\in\mathcal{E}^{\pi}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
Thus a function $f\in\mathcal{E}^{\pi}$ is only determined by the values $f(1),f(2),f(3),\dots$.
Remark 2.
The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}^{\pi}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty$
2.4 Ramanujan’s method of summation
Let $f$ and $g$ two functions verifying the hypothesis of theorem 1. By Carlson’s theorem we know that if $f\in\mathcal{E}^{\pi}$ and $g\in\mathcal{E}^{\pi}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
f(x+it)f(xit)=g(x+it)g(xit),{f(x+it)-f(x-it)}={g(x+it)-g(x-it)},
thus, by (4), we have $C(f)=C(g).$ This leads to the following definition.
Definition 2.
For $f\in\mathcal{E}^{\pi}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
It is clear that $\sum_{n\geq 1}^{\mathcal{R}}f(n)$ depends linearly on $f$.
Remark 3.
Let $f\in\mathcal{E}^{\pi}$ then by (3), we have for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\!\int_{1}^{n}f(u)\ du+\frac{f(n)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (7)
If $\lim_{n\to+\infty}f(n+it)=0$, for every t\ $\in\mathbb{R}, t$hen, by dominated convergence, we get
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,=0,$
and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case we have by (7)
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\sum_{n=1}^{+\infty}f(n)-\!\!\int_{1}^{+\infty}\!\!f(x)\ dx.$
Example 2.
For $f:x\mapsto\frac{1}{x^{s}}$, with $\Re(s)>1$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\sum_{n=1}^{+\infty}\frac{1}{n^{s}}-\!\!\int_{1}^{+\infty}\!\!\frac{1}{x^{s}}\ dx,$
thus, for $\Re(s)>1,$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\zeta(s)-\frac{1}{s-1}.$ (8)
2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of partial sums of this series . In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
${\varphi(x)-\varphi(x-1)=f(x)}\ \text{ with }\varphi(0)=0.$
The following theorem clarifies this point.
Theorem 3.
Let $f\in\mathcal{E}^{\pi}$. There is a unique function $\varphi_{f}\in\mathcal{E}^{\pi}$, such that
[…]
Proof.
a) Existence. Since $f\in\mathcal{E}^{\pi}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
$\varphi_{f}(z)=\int_{1}^{z}f(u)\ du+\frac{f(z)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(z+it)-f(z-it)}{e^{2\pi t}-1}\ dt,$ (10)
is also in $\mathcal{E}^{\pi}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
$\varphi_{f}(n+1)-\varphi_{f}(n)\,=f(n+1),\text{ for }n=1,2,3,\dots.$
Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}^{\pi}$, then by Carlson’s theorem, we get
$\varphi_{f}(z+1)-\varphi_{f}(z)=f(z+1),\quad\!\!\text{ for all }z\text{ such that }\Re(z)>0.\$(11)
The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (6) and (10).
b) Uniqueness. If $\varphi\in\mathcal{E}^{\pi}$ is such that
$\varphi(x+1)-\varphi(x)=f(x+1)\ \text{ for }x>0,\ \ \text{ with }\varphi(1)=f(1),$
then, for every integer $n\geq 1$, we get
$\varphi(n)-f(1)=\varphi(n)-\varphi(1)\;=\sum_{k=1}^{n-1}\big(\varphi(k+1)-\varphi(k)\big)=\sum_{k=1}^{n-1}f(k+1),$
thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$. Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}^{\pi}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
Remark 4.
The preceding proof asserts that if $f\in\mathcal{E}^{\pi}$ then
$\varphi_{f}(n)=\sum_{k=1}^{n}f(k)\text{ for every integer }n\geq 1.$
The function $\varphi_{f}$ given by (10) is the unique function in $\mathcal{E}^{\pi}$ wich interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
$\varphi_{f}(z)=\varphi_{f}(z+1)-f(z+1),$
and since the right-hand side defines an analytic function in the half-plane $\{\Re(z+1)>0\}$, this relation gives an analytic continuation of $\varphi_{f}$ in the half-plane $\{\Re(z)>-1\}.$ Let’s denote again by $\varphi_{f}$ this continuation, then the relation $\varphi_{f}(1)=f(1)$} provides the condition required by Ramanujan
$\varphi_{f}(0)=\varphi_{f}(1)-f(1)=0.$}
Example 3.
Let $f:z\mapsto 1/z$, then $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
$\Psi(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}.$
This function is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]) we get for $\Re(z)>0$
$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}$ (12)
Then it is easily verified that $\Psi\in\mathcal{E}^{\pi}$ and that
$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}$
Since $\Psi(1)=-\gamma$, then, by the previous theorem, we get for $\Re(z)>0$
$\varphi_{f}(z)=\Psi(z+1)+\gamma.$
A very useful formula for Ramanujan’s summation is provided by the following theorem.
Theorem 4.
Let $f\in\mathcal{E}^{\pi}$, and let $\varphi\in\mathcal{E}^{\pi}$ be a solution of
$\varphi(x+1)-\varphi(x)$ $=$ $f(x+1),\ \text{ for all }\ x>0,$
$\varphi(1)$ $=$ $f(1).$
Then we have
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\int_{0}^{1}\varphi(x)dx.$
Proof.
By theorem 3, we have $\varphi=\varphi_{f}$. Let $F(z)=\int_{1}^{z}f(u)\ du$, then integration of (10) gives
$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$ (13)
where
$D(f)=\int_{1}^{2}\Big(\int_{0}^{+\infty}\frac{f(x+it)-f(x-it)}{e^{2\pi t}-1}\,\ dt\Big)\ dx.$
By Fubini’s theorem we have
$D(f)\;\;=\int_{0}^{+\infty}{\Big(\int_{1}^{2}{f(x+it)-f(x-it)}\ dx\Big)}\frac{1}{e^{2\pi t}-1}dt,$
thus
$D(f)=\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt-\!\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt.$
If we apply Plana’s formula to the function $F$, with $n=2$, then we get
$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)dt$ $=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt$
$-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt,$
which is
$iD(f)=-\frac{F(2)}{2}+\int_{1}^{2}F(t)\ dt.$
Then (13) becomes
$\int_{1}^{2}\varphi_{f}(x)\ dx$ $=$ $\frac{1}{2}\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n)+\frac{F(2)}{2}$
$=$ $\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n).$
from which we deduce that
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\int_{0}^{1}\big(\varphi_{f}(x+1)-f(x+1)\big)\ dx=\int_{0}^{1}\varphi_{f}(x)\ dx.$
Corollary 1.
Let $f\in\mathcal{E}^{\pi}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}^{\pi}$ and we have the shift formula
$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$
Proof.
Let $g:x\mapsto f(x+1)$, then it is easily verified, by theorem 3, that we have for $x>0$
$\varphi_{g}(x)=\varphi_{f}(x+1)-f(1).$
Thus we have
$\int_{0}^{1}\varphi_{g}(x)dx=\int_{0}^{1}\varphi_{f}(x+1)dx-f(1)=\int_{0}^{1}\varphi_{f}(x)dx+\int_{0}^{1}f(x+1)dx-f(1),$
and the conclusion follows from theorem 4. ∎
Example 4.
Let $f:x\mapsto 1/x$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\int_{0}^{1}(\Psi(x+1)+\gamma)dx=\int_{0}^{1}\frac{\Gamma^{\prime}}{\Gamma}(x+1)dx+\gamma,$
thus we verify again that
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\gamma,$ (14)
Observe that, by corollary 1, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}=\gamma-1+\log(2).$
Example 5.
Since, fro any $x>0$, we have
$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1)$
and $\log(\Gamma(2))=\log(1)$, we deduce that $\varphi_{\log}(x)=\log(\Gamma(x+1))$. Now if we use the Raabe’s formula, that is :
$\int_{a}^{a+1}\log(\Gamma(x))dx=a\log(a)-a+\log(\sqrt{2\pi}),$
then we get
$\sum_{n\geq 1}^{\mathcal{R}}\log(n)=\log(\sqrt{2\pi})-1.$
Example 6.
Let $f(x)=\Psi(x)$ and $\varphi(x)=x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
$\varphi(x+1)-\varphi(x)=\Psi(x+1)=f(x+1).$
Thus, by the previous theorem, we get
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=\int_{0}^{1}x\Psi(x)dx+\frac{1}{2}=-\int_{0}^{1}\log(\Gamma(x))dx+\frac{1}{2},$
then by Raabe’s formula
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$
Example 7.
We can evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for $k=1,2,3,\dots$ by means of Bernoulli numbers. Let the Bernoulli polynomials $B_{n}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
$B^{\prime}_{k}(x)=nB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)dx=0.$
The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. These polynomials verify (cf. [G-R])
$B_{k+1}\big((x+1)+1\big)-B_{k+1}(x+1)=(k+1)(x+1)^{k}.$
For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. Let the function $f:x\mapsto x^{k}$ with an integer $k\geq 1$, then by theorem 3, we obtain
$\varphi_{f}(x)=\frac{B_{k+1}(x+1)}{k+1}-\frac{B_{k+1}}{k+1}.$
Since
$\int_{0}^{1}\frac{B_{k+1}(x+1)}{k+1}dx\;\;=\int_{0}^{1}\frac{B_{k+1}(x)}{k+1}dx\,+\int_{0}^{1}x^{k}dx=\frac{1}{k+1},$
then theorem 4 gives
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}=\frac{1-B_{k+1}}{k+1}.$ (15)
Remark 5.
Relations (14) and (15) can be written in terms of the Zeta function :
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}$ $=$ $\lim_{s\to 1}\Big(\zeta(s)-\frac{1}{s-1}\Big),$
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ $=$ $\zeta(-k)-\frac{1}{-k-1}\text{ for every integer }k\geq 1.$
Note also that
$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1,$
and by (9) we have in the half-plane $\{\Re(s)>1\}$
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}.$ (16)
All these formulae are special cases the following general result.
Theorem 5.
The function
$s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;$
is analytic on $\mathbb{C}$ and we have for every $s\in\mathbb{C}\setminus\{1\}$
$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}$ (17)
Proof.
First, it is easily seen that, for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}^{\alpha}$, for all $\alpha>0.$ Since, if $s=\sigma+i\tau$ ($\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$), we can write
$\Big|\frac{1}{z^{s}}\Big|=\Big|\frac{1}{(x+it)^{s}}\Big|=\Big|e^{-s\log(x+it)}\Big|=e^{-{\sigma}\log\sqrt{x^{2}+t^{2}}+\tau\arctan(\frac{t}{x})}\leq e^{|\tau|\pi/2}(x^{2}+t^{2})^{-\sigma/2}$
where $z=x+it$, with $x>\varepsilon>0$ and $t\in\mathbb{R}$. Thus the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined and, by (6), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\;\;\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt\,$
$\frac{1}{2}+2\int_{0}^{+\infty}\frac{\sin(s\arctan(t))}{(1+t^{2})^{s/2}}\frac{1}{e^{2\pi t}-1}dt.$
This last integral depends analyticaly of $s\in\mathbb{C}$ so the function $s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is analytic on $\mathbb{C}$. Since the function $s\mapsto\zeta(s)-\frac{1}{s-1}$ is analytic on $\mathbb{C}\setminus\{1\}$ then, by analytic continuation of (16), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}\text{ for any }s\in\mathbb{C}\setminus\{1\}.$
3 An expansion of Zeta in the critical strip
We are now going to use expression (17) to obtain, in the strip $\{0<\Re(s)<1\}$, a polynomial expansion of Zeta
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s).$
As we have seen in the Introduction the polynomials $P_{m}$ are related to the Laguerre polynomials by Mellin transform. We now give another approach to these polynomials.
3.1 The sequence of polynomials $(P_{m})$
We start with the observation that for $s\in\mathbb{C}$, and for every $n>0$, we have
$\frac{1}{n^{s}}=\frac{1}{n+1}G_{s}(\frac{n-1}{n+1}),$ (18)
where $G_{s}$ is the function defined on $D(0,1)=\{z\in\mathbb{C};|z|<1\}$ by
$G_{s}:z\mapsto\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}$ (19)
Since $\Re\big({\frac{1-z}{1+z}}\big)>0$ for $z\in D(0,1)$, we see that the function $G_{s}$ is analytic on $D(0,1)$. Let $P_{m}(s)$ denote the coefficients of its Taylor expansion
$\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)z^{m}\ \text{ for }\ \ \ z\in D(0,1)$ (20)
Let $z=\frac{x}{1+x}$, with $x\in]0,1[$, then, using (20), we get
$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$ (21)
where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s(s+1)\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ We use now the well-known equivalence related to Euler summation (cf.[H1])
$\sum_{m\geq 0}b_{m}(\frac{x}{1+x})^{m+1}=\sum_{k\geq 0}a_{k}x^{k+1}\ \Leftrightarrow\ b_{m}=\sum_{k=0}^{m}\binom{m}{k}\ a_{k},$
which gives, for $P_{m}(s),$ the following explicit expression
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{(s)_{k}}{k!}.$ (22)
We also have an integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
$\frac{(s)_{k}}{k!}=\frac{1}{\Gamma(s)\Gamma(1-s)}\int_{0}^{1}t^{s-1+k}(1-t)^{-s}dt,$
thus, by (22), we get for $0<\Re(s)<1$
$P_{m}(s)=2\frac{\sin(\pi s)}{\pi}\int_{0}^{1}t^{s-1}(1-t)^{-s}(1-2t)^{m}dt.$ (23)
This integral expression allows us to obtain the following lemma which will be useful later.
Lemma 1.
For $0<\Re(s)<1$, we have, for every $m\geq 1$
$|P_{m}(s)|\leq e^{\pi|Im(s)|}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big).$
Proof.
Using (23) we get
$|P_{m}(s)|$ $\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{0}^{1}t^{\Re(s)-1}(1-t)^{-\Re(s)}|1-2t|^{m}dt,$
$\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}du,$
$\leq$ $e^{\pi|Im(s)|}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}du.$
Let $a=\Re(s)$, this last integral is :
$I_{a}=\int_{0}^{1}(1-u)^{a-1}(1+u)^{-a}u^{m}du+\int_{-1}^{0}(1-u)^{a-1}(1+u)^{-a}(-u)^{m}du.$
Since $0<a<1$, we have
$I_{a}\leq\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{-1}^{0}(1+u)^{-a}(-u)^{m}du$ $=$ $\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{0}^{1}(1-u)^{-a}u^{m}du.$
Now if $0<\alpha<1$ we have, for every $x>0$
$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{x}du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-xt}dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-xt}dt=\Gamma(\alpha)x^{-\alpha},$
(the last inequality is justified by $1\leq\frac{t}{1-e^{-t}}\leq{e^{t}}$, for every $t>0$). Finally we obtain
$I_{a}\leq\frac{\Gamma(a)}{m^{a}}+\frac{\Gamma(1-a)}{m^{1-a}}.$
Remark 6.
It is easily verified (cf. [G-R]) that, for $m\geq 1$, we have $P_{m}(s)=2F(-m,s,1,2),$ where $F$ is the Gauss hypergeometric function
$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!},$
By classical properties of hypergeometric function (cf. [G-R]), we get the relations
$P_{m}(1-s)=(-1)^{m}P_{m}(s),$ (24)
$sP_{m}(s+1)-(s-1)P_{m}(s-1)=(2m+1)P_{m}(s).$ (25)
3.2 Orthogonality relations
Let $(\varphi_{m})_{m\geq 0}$ the sequence of functions
$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x).$
It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),dx)$. We verify as in the introduction that, by Mellin transform, we have for $\Re(s)>0$
$\mathcal{M}\varphi_{m}(s)=\int_{0}^{+\infty}\varphi_{m}(x)x^{s-1}\ dx=\frac{1}{\sqrt{2}}{\Gamma(s)}P_{m}(s).$ (26)
Thus the Mellin-Parseval formula (cf. [T]), gives
$\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(\varphi_{n})(\frac{1}{2}+it)\ dt=\int_{0}^{+\infty}\overline{\varphi_{m}(x)}\varphi_{n}(x)dx=\delta_{(m,n)},$
with $\delta$ the usual Kronecker symbol. That is
$\frac{1}{4\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)}\overline{P_{m}(\frac{1}{2}+it)}\ \Gamma(\frac{1}{2}+it)P_{n}(\frac{1}{2}+it)dt=\delta_{(m,n)}.$
Since
$\Gamma(\frac{1}{2}+it)\overline{\Gamma(\frac{1}{2}+it)}=\frac{\pi}{\cosh\pi t},$
then
$\frac{1}{4}\int_{-\infty}^{+\infty}\overline{P_{m}(\frac{1}{2}+it)}\ P_{n}(\frac{1}{2}+it)\frac{dt}{\cosh\pi t}=\delta_{(m,n)}.$
This proves that the sequence $(Q_{m})_{m\geq 0}$ of polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}\ \frac{(\frac{1}{2}+it)_{k}}{k!}$ (27)
is an orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. It is also an orthonormal basis of this space since for $|a|<\pi$, we have
$\int_{-\infty}^{+\infty}e^{a|t|}\frac{dt}{\cosh(\pi t)}<+\infty$
By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
For the first values of $m$ we get
$Q_{0}(t)$ $=$ $1,$
$Q_{1}(t)$ $=$ $i(-2t),$
$Q_{2}(t)$ $=$ $-2t^{2}+\frac{1}{2},$
$Q_{3}(t)$ $=$ $i(\frac{4}{3}t^{3}-\frac{5}{3}t),$
$Q_{4}(t)$ $=$ $\frac{2}{3}t^{4}-\frac{7}{3}t^{2}+\frac{3}{8}.$
Remark 7.
By (24) and (25) we have, for every $m\geq 0$, the relations
$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
$\big(\frac{1}{2}+it\big)Q_{m}(t-i)+\big(\frac{1}{2}-it\big)Q_{m}(t+i)=(2m+1)Q_{m}(t).$
3.3 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
We now apply Ramanujan’s summation on both sides of (18). Using (20), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}G_{s}(\frac{n-1}{n+1})=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
The following theorem says that, for $0<\Re(s)<1$, these last two sums can be inverted.
Theorem 6.
For $0<\Re(s)<1$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
And we obtain the expansion
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}P_{m}(s).$ (28)
Proof.
By definition 2, we can write
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=c_{m}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt$
with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. We will now prove that the series
$\sigma(s)=\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
is convergent for every $0<\Re(s)<1$, and that its sum is equal to
$\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
Since $c_{0}P_{0}(s)=1/2$, then, by definition 2, the series $\sigma(s)$ is also
$\sigma(s)=\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$ (29)
To prove the convergence of this series and invert signs $\sum$ and $\int$ we use the following lemma.
Lemma 2.
For $0<\Re(s)<1$, the series
$\sum_{m\geq 0}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|$ (30)
is convergent and
$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}dt<+\infty.$ (31)
Proof of the lemma.
We have
$\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big|\leq 2\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}\Big|=\frac{2}{\sqrt{t^{2}+4}}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}$
Using lemma 1, we see that the terms of the series (30) are, for $m\geq 1$, dominated by the terms
$\frac{4}{\sqrt{t^{2}+4}}e^{\pi|Im(s)|}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
Consequently, the series (30) is convergent, and we have
$\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \leq M(t)$
with
$M(t)=\frac{4t}{4+t^{2}}+e^{\pi|Im(s)|}\frac{4}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
It remains to prove that
$\int_{0}^{+\infty}M(t)\frac{1}{e^{2\pi t}-1}dt<+\infty.$
This follows from the convergence of the integral
$\int_{0}^{+\infty}\frac{4t}{4+t^{2}}\frac{1}{e^{2\pi t}-1}dt,$
and the fact that, for $0<\alpha<1$, we have
$\int_{0}^{+\infty}\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\frac{1}{e^{2\pi t}-1}dt<+\infty,$
as a consequence of the following inequalities
$\frac{1}{\sqrt{t^{2}+4}}\sum_{m=2}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\leq\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\leq\frac{t}{{t^{2}+4}}\Big(\frac{1}{1-\frac{t}{\sqrt{t^{2}+4}}}\Big)\leq\frac{t^{2}}{4}.$
End of the proof of the lemma.
By the preceding lemma the series $\sigma(s)$ is convergent for every $0<\Re(s)<1$, and we get
$\sum_{m=0}^{+\infty}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\sum_{m=0}^{+\infty}\big(\frac{it}{2+it}\big)^{m}P_{m}(s)-\frac{1}{2-it}\sum_{m=0}^{+\infty}\big(\frac{-it}{2-it}\big)^{m}P_{m}(s)\Big)\frac{1}{e^{2\pi t}-1}\ dt.$
Using (20) with $z=\frac{\pm it}{2\pm it}$, this last expression is equal to
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt,$
which is by definition 2, also equal to
$\frac{1}{2}+\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}-\frac{1}{2}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
The relation (28) follows immediately from (17).
3.4 Evaluation of the sum $z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
If we apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+y_{m}$
with
$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx=\int_{1}^{2}\Big(1-\frac{2}{x}\Big)^{m}\frac{1}{x}dx.$
Now the binomial expansion
$\Big(1-\frac{2}{x}\Big)^{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\frac{1}{x^{k}}$
gives
$z_{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k+1}}-(-1)^{m}+\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}dx.$
Thus we obtain $z_{m}=x_{m}+y_{m}$, where $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\Big(\zeta(k+1)-\frac{1}{k}\Big)-(-1)^{m},$ (32)
$y_{m}$ $=$ $\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$
For the first values of $m$ we get
$x_{0}$ $=$ $\gamma-1,$
$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3,$
$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
Theorem 7.
For $0<\Re(s)<1$, we have
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}P_{m}(s),$ (33)
where
$z_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\frac{1}{k}.$
Proof.
This follows from the expansion (28) of theorem 6, and the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1}{k}.$
3.5 An expansion of Zeta
The previous theorem will enable us to get an expansion of Zeta in the strip $\{0<\Re(s)<1\}$. Indeed, the polar part 1/(s1)1/(s-1) of (33) has the simple expansion given by the following lemma.
Lemma 3.
For $0<\Re(s)<1$ we have
$\frac{1}{s-1}=-\sum_{m=0}^{+\infty}y_{m}P_{m}(s).$ (34)
Proof.
For $0<\Re(s)<1$, we write
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\frac{1}{x+1}dx.$ (35)
This interchanging of signs $\sum$ and $\int$ is justified by
$\sum_{m=0}^{+\infty}\int_{0}^{1}|\frac{x-1}{x+1}|^{m}\ |P_{m}(s)|\frac{1}{x+1}dx\leq\sum_{m=0}^{+\infty}\int_{0}^{1}\frac{1}{(x+1)^{m+1}}dx\ |P_{m}(s)|\leq\sum_{m=0}^{+\infty}\frac{1}{m}|P_{m}(s)|,$
and we immediately check, with the lemma 1, that the latter series is convergent.
The relations (18) and (20) allows us to write for every $x>0$
$\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)=\frac{1}{x^{s}},$
and by (35) we have
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\int_{0}^{1}\frac{1}{x^{s}}dx=-\frac{1}{s-1}.$
Theorem 8.
For $0<\Re(s)<1$, we have
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$ (36)
with
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$ (37)
and $r_{m}$ are the rational numbers
$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}$
Proof.
This is an immediate consequence of the preceding lemma and theorem 7 and the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-2)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
3.6 An orthonormal expansion
Let $X(0)=\gamma-1$ and for $k\geq 1$
$X(k)=\zeta(k+1)-1-\frac{1}{k}.$
Then formula (32) can be written as a binomial transform
$x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$
Let
$f_{0}(t)=\frac{1}{e^{t}-1}-\frac{1}{t},$
then we have for $k\geq 0$
$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)dt$
By definition of Laguerre polynomials, we obtain for the coefficients $x_{m}$ the integral expression
$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$ (38)
Theorem 9.
We have for any integer $m\geq 0$
$x_{m}=\frac{1}{2}\int_{-\infty}^{+\infty}\overline{Q_{m}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$ (39)
Proof.
By Mellin-Parseval formula (cf. [T]), we have
$\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt$
Since (cf. [T]) we have $\mathcal{M}f(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$, then, using (26), we get
$x_{m}=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt,$
since $Q_{m}(t)=\frac{1}{2}P_{m}(\frac{1}{2}+it)$, this is the integral expression (39).
The previous theorem means that the pointwise convergent expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
is also an orthonormal expansion in the space $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$.
Then we have
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=4\sum_{m=0}^{+\infty}x_{m}^{2}.$ (40)
We give next a formal evaluation of this last sum using Ramanujan’s summation.
3.7 A generating function
Let $\mathcal{L}$ the Laplace transform
$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx=(e^{-xz},f).$
By Paley-Wiener theorem (cf. [Ho], [P]) , the transformation $\frac{1}{\sqrt{2\pi}}\mathcal{L}$ maps isometrically the space $L^{2}(]0,+\infty[,dx)$ to the Hardy space $H^{2}(\mathbb{C}_{+})$ of analytic functions $g$ in the half-plane $\mathbb{C}_{+}=\{z|\Re(z>0\}$, such that : there exists $M_{g}>0$ with
$\int_{\mathbb{R}}|g(x+iy)|^{2}\ dy\leq M_{g}\text{ for all }x>0.$
Each such $g$ has non-tangential limits $g(iy)=\lim_{x\to 0}g(x+iy)$ at almost every point of the imaginary axis. The space $H^{2}(\mathbb{C}_{+})$ is an Hilbert space with the inner product
$(g,h)_{H^{2}}=\int_{\mathbb{R}}\overline{g(iy)}{h(iy)}dy,$
and the functions $\psi_{m}$ $(m\geq 0)$, form an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. If $f\in L^{2}(]0,+\infty[,dx)$. From
$e^{-xz}=\sqrt{2}\ \sum_{k=0}^{+\infty}\varphi_{k}(x)\psi_{k}(z).$
we get
$\frac{1}{\sqrt{2\pi}}\mathcal{L}\varphi_{m}(z)=\frac{1}{\sqrt{2\pi}}(e^{-xz},\varphi_{m})=\frac{1}{\sqrt{\pi}}\psi_{m}(z)$
The system $(\frac{1}{\sqrt{\pi}}\psi_{m})_{m\geq 0}$ is an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. From the expansion
$f_{0}=\sum_{m\geq 0}(\varphi_{m},f_{0})\varphi_{m}$
we get the expansion
$\frac{1}{\sqrt{2\pi}}\mathcal{L}f_{0}(z)=\sum_{m\geq 0}(\varphi_{m},f_{0})\frac{1}{\sqrt{\pi}}\psi_{m}(z).$
and, by (38) we deduce
$\mathcal{L}f_{0}(z)=2\sum_{m\geq 0}x_{m}\psi_{m}(z).$
Since it is well known (cf. [G-R]) that
$\int_{0}^{+\infty}e^{-xz}f_{0}(x)dx=\log(z)-\Psi(z)-\frac{1}{z}$
then we get
$2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(z)=\log(z)-\Psi(z)-\frac{1}{z}.$ (41)
Equivalently this gives the generating function
$\sum_{m=0}^{+\infty}x_{m}u^{m}=\frac{1}{1-u}\Big(\log(\frac{1+u}{1-u})-\Psi(\frac{1+u}{1-u})\Big)-\frac{1}{1+u}.$ (42)
Remark 8.
Now, if we apply (41) with $z=1,2,3,\dots$ and take Ramanujan’s summation, we get
$2\sum_{m=0}^{+\infty}x_{m}\sum_{n\geq 1}^{\mathcal{R}}\psi_{m}(n)=2\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}x_{m}\psi_{m}(n)=\sum_{n\geq 1}^{\mathcal{R}}\log(n)-\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)-\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}.$
Then, with the results of examples 4-6, this relation gives
$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{n})=\log(2\pi)-\gamma-\frac{3}{2}.$
Since $y_{m}=\int_{0}^{1}\psi_{m}(x)dx$ then by (41) we have
$2\sum_{m=0}^{+\infty}x_{m}y_{n}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)dx=\int_{0}^{1}(\log(x)-\Psi(1+x))dx=-1.$
Finally we get
$\sum_{m=0}^{+\infty}x_{m}^{2}=\frac{1}{2}\log(2\pi)-\frac{1}{2}\gamma-\frac{1}{4}.$
Then (40) gives us the relation
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=2\log(2\pi)-2\gamma-1,$ (43)