### v4

# arXiv:2512.11405v4[math.NT] 01 Mar 2026

Abstract
We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann’s zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$. The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
1 Introduction
Some polynomial expansions of the Riemann’s zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to the zeta function. For example
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}c_{m}\binom{s-1}{m}$,
with
$c_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-1)^{k}\big(\zeta(k+1)-\frac{1}{k}\big).$
Others (cf.[E-G]) are linked to the well known expansion
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}\gamma_{m}\frac{(s-1)^{m}}{m!}$
where $\gamma_{m}$ are the Stieltjes constants.
The polynomial expansion of the Riemann’s zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
$L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\frac{x^{k}}{k!}.$
The generating function of the Laguerre polynomials (cf. [Leb]) gives for $x\in[0,+\infty[$
$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\big(\frac{z-1}{z+1}\big)^{m}\frac{1}{1+z}.$
By Mellin transform of the function $x\mapsto e^{-xz}$, we obtain for $\Re(z)>0$ and $\Re(s)>0$
$z^{-s}=\sum_{m=0}^{+\infty}\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}P_{m}(s),$
where $P_{m}$ is a polynomial defined by
${\Gamma(s)}P_{m}(s)=2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$
Take $z=1,2,3,\dots$, then we get an infinite sequence of relations that it’s tempting to sum up, to obtain
$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\frac{1}{1+n}\big(\frac{n-1}{n+1}\big)^{m}\Big)P_{m}(s).$
We prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
We give in section 2, a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1},$
then, after an explicit evaluation of $z_{m}$, we obtain, for $0<\Re(s)<1$, the following expansion
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s),$
with
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$
where $r_{m}$ is given by
$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
We prove that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. This is a consequence of the fact that the polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
form a complete orthonormal system in $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. The expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
is an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Also by a formal calculation, we get
$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$
2 Ramanujan’s summation
2.1 Ramanujan’s constant of a series
Let a function $f\in C^{1}(]0,+\infty[)$, then for every integer $n\geq 1$, we have the first order Euler-MacLaurin formula
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}b(x)f^{\prime}(x)\ dx$}, (1)
where
b(x)=x[x]12 and [x] denote the integer part of x.b(x)=x-[x]-$\frac${1}{2}\ $\text${ and }\ [x]\ $\text${ denote the integer part of }\ x.
Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. Then the previous formula gives
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+C(f)+\frac{1}{2}f(n)-\int_{n}^{+\infty}b(x)f^{\prime}(x)\ dx$}, (2)
 $\text{ with }\ \ C(f)=\frac{1}{2}f(1)+\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx.$
The constant $C(f)$ is called the Ramanujan constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
Clearly $C(f)$ depends linearly of $f$. And if the series $\sum_{n\geq 1}f(n)$ is convergent, then by (2) we have
$C(f)=\sum_{k=1}^{+\infty}f(k)-\int_{1}^{+\infty}f(x)\ dx.$
We observe that $C(f)$ is also well defined for some divergent series. For example, we have immediately $C(1)=1/2,$ and applying (2) to the function $f:x\mapsto 1/x$, we have for any integer $n\geq 1$
$\sum_{k=1}^{n}\frac{1}{k}=\log n+C(f)+\frac{1}{2n}+\int_{n}^{+\infty}b(x)\frac{1}{x^{2}}\ dx,$
and we see that the Ramanujan constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
$\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
2.2 Plana’s formula
The following summation formula was discovered by Plana in 1820 (and also given by Abel in 1823) (cf. [Lin]).
Theorem 1.
Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $M>0$ with
$\left|f(x+it)\right|\leq M\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and }t\in\mathbb{R}.$
Then we have Plana’s formula : for any integer $n\geq 1$,
$\sum_{k=1}^{n}f(k)$ $=\frac{f(1)+f(n)}{2}+\!\int_{1}^{n}f(u)\ du$
$+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt-i\!\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (3)
If we now compare Euler-MacLaurin formula (1) and Plana’s formula (3), then we get, for $n\geq 1$,
$\int_{1}^{n}b(x){f^{\prime}(x)}\ dx=i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$
Then if the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\ dt=0,$
we have
$\int_{1}^{+\infty}b(t){f^{\prime}(t)}\ dt=i\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt.$
Thus, in this case the Ramanujan constant $C(f)$, previously defined in (2), is also
$C(f)=\frac{1}{2}f(1)+i\int_{0}^{+\infty}\frac{{f(1+it)-f(1-it)}}{e^{2\pi t}-1}\,\ dt.$ (4)
Remark 1.
Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$. Then the Ramanujan constants $C(f)$ and $C(g)$, defined by (4), need not be the same. For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
$C(g)=i\int_{0}^{+\infty}\frac{\sin\big(\pi(1+it)\big)-\sin\big(\pi(1-it)\big)}{e^{2\pi t}-1}\,\ dt=\int_{0}^{+\infty}\frac{e^{\pi t}-e^{-\pi t}}{e^{2\pi t}-1}\,\ dt\;\;=\;\int_{0}^{+\infty}e^{-\pi t}\ dt=\frac{1}{\pi}.$
To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that the Ramanujan constant, defined by (4), only depends on the values $f(n)$ for all integers $n\geq 1$}.
2.3 The space $\mathcal{E}$
Definition 1.
Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there are some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
for every ${\varepsilon}>0$, there is a constant $C_{\varepsilon}$ such that
$\left|f(z)\right|\;\;\leq\;\;C_{\varepsilon}\ e^{b\,\Re(z)}\ e^{a\,|\Im(z)|},\ \text{ if }\ \Re(z)>{\varepsilon}.$ (5)
Example 1.
Let $f_{s}\;:\;z\mapsto e^{sz}$, with $s=\sigma+i\tau$, $\sigma\in\mathbb{R}$ and $\tau\in\mathbb{R}$ . For $\Re(z)>0$, we have
$|f_{s}(z)|\;\;=\,\;e^{\sigma\Re(z)-\tau\Im(z)}.$
Thus $f_{s}\in\mathcal{E}$ if and only if $\pi<\Im(s)<\pi$.
An important theorem for the rest of this section is the following.
Theorem 2.
Carlson’s theorem (cf. [H2])
If $f\in\mathcal{E}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
Thus a function $f\in\mathcal{E}$ is only determined by the values $f(1),f(2),f(3),\dots$.
Remark 2.
The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty.$
2.4 Ramanujan’s method of summation
Let $f$ and $g$ two functions verifying the hypothesis of theorem 1. By Carlson’s theorem we know that if $f\in\mathcal{E}$ and $g\in\mathcal{E}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
f(x+it)f(xit)=g(x+it)g(xit),{f(x+it)-f(x-it)}={g(x+it)-g(x-it)},
thus, by (4), we have $C(f)=C(g).$ This leads to the following definition.
Definition 2.
For $f\in\mathcal{E}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=C(f)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
It is clear that $\sum_{n\geq 1}^{\mathcal{R}}f(n)$ depends linearly on $f$.
Remark 3.
Let $f\in\mathcal{E}$ then by Plana’s formula (3), we have for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\!\int_{1}^{n}f(u)\ du+\frac{f(n)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (7)
If $\lim_{n\to+\infty}f(n+it)=0$, for every t\ $\in\mathbb{R}, t$hen, by dominated convergence, we get
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,=0,$
and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case (7) gives the relation
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\sum_{n=1}^{+\infty}f(n)-\!\!\int_{1}^{+\infty}\!\!f(x)\ dx.$
Example 2.
Let $f:x\mapsto\frac{1}{x^{s}}$, then for $\Re(s)>1$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\sum_{n=1}^{+\infty}\frac{1}{n^{s}}-\!\!\int_{1}^{+\infty}\!\!\frac{1}{x^{s}}\ dx,$
thus, for $\Re(s)>1$, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\zeta(s)-\frac{1}{s-1}.$ (8)
2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler formula than (6) can be obtained by means of an interpolation function of the partial sums of this series. In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
${\varphi(x)-\varphi(x-1)=f(x)}\ \text{ with }\varphi(0)=0.$
The following theorem clarifies this point.
Theorem 3.
Let $f\in\mathcal{E}$. There exists a unique function $\varphi_{f}\in\mathcal{E}$, such that
[…]
Proof.
a) Existence. Since $f\in\mathcal{E}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
$\varphi_{f}(z)=\int_{1}^{z}f(u)\ du+\frac{f(z)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(z+it)-f(z-it)}{e^{2\pi t}-1}\ dt,$ (10)
is also in $\mathcal{E}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
$\varphi_{f}(n+1)-\varphi_{f}(n)\,=f(n+1),\text{ for }n=1,2,3,\dots.$
Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}$, then by Carlson’s theorem, we get
$\varphi_{f}(z+1)-\varphi_{f}(z)=f(z+1),\quad\!\!\text{ for all }z\text{ such that }\Re(z)>0.\$(11)
The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (10) and definition (6).
b) Uniqueness. If $\varphi\in\mathcal{E}$ is such that
$\varphi(x+1)-\varphi(x)=f(x+1)\ \text{ for }x>0,\ \ \text{ with }\varphi(1)=f(1),$
then, for every integer $n\geq 1$, we get
$\varphi(n)-f(1)=\varphi(n)-\varphi(1)\;=\sum_{k=1}^{n-1}\big(\varphi(k+1)-\varphi(k)\big)=\sum_{k=1}^{n-1}f(k+1),$
thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$. Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
Remark 4.
The preceding proof asserts that if $f\in\mathcal{E}$ then the function $\varphi_{f}$ given by (10) satisfies
$\varphi_{f}(n)=\sum_{k=1}^{n}f(k)\text{ for every integer }n\geq 1.$
Thus this function $\varphi_{f}$ is the unique function in $\mathcal{E}$ which interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
$\varphi_{f}(z)=\varphi_{f}(z+1)-f(z+1),$
and since the right-hand side defines an analytic function in the half-plane $\{\Re(z+1)>0\}$, this relation gives an analytic continuation of $\varphi_{f}$ in the half-plane $\{\Re(z)>-1\}.$ Let’s denote again by $\varphi_{f}$ this continuation, then the relation $\varphi_{f}(1)=f(1)$} provides the condition required by Ramanujan
$\varphi_{f}(0)=\varphi_{f}(1)-f(1)=0.$}
Example 3.
Let the function $f:z\mapsto 1/z$. We will ses that $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
$\Psi(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}.$
The function $\Psi$ is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]), we get for $\Re(z)>0$
$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}.$
Thus it is easily verified that $\Psi\in\mathcal{E}$ and that
$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}.$
Since $\Psi(1)=-\gamma$, then, by the previous theorem, we get for $\Re(z)>0$
$\varphi_{f}(z)=\Psi(z+1)+\gamma.$
Theorem 4.
Let $f\in\mathcal{E}$, and let $\varphi\in\mathcal{E}$ be a solution of
$\varphi(x+1)-\varphi(x)=f(x+1),\ \text{ for all }\ x>0,\ \text{ and }\ \varphi(1)=f(1).$
Then we have
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\int_{0}^{1}\varphi(x)\ dx.$
Proof.
By theorem 3, we have $\varphi=\varphi_{f}$. Let $F(z)=\int_{1}^{z}f(u)\ du$, then integration of (10) gives
$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$
where
$D(f)=\int_{1}^{2}\Big(\int_{0}^{+\infty}\frac{f(x+it)-f(x-it)}{e^{2\pi t}-1}\,\ dt\Big)\ dx.$
By Fubini’s theorem we have
$D(f)\;\;=\int_{0}^{+\infty}{\Big(\int_{1}^{2}{f(x+it)-f(x-it)}\ dx\Big)}\frac{1}{e^{2\pi t}-1}\ dt,$
thus
$D(f)=\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}\ dt-\!\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}\ dt.$
If we apply Plana’s formula (3) to the function $F$, in the case $n=2$, then we get
$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)\ dt=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}\ dt,$
that is
$iD(f)=-\frac{F(2)}{2}+\int_{1}^{2}F(t)\ dt.$
Then we get
$\int_{1}^{2}\varphi_{f}(x)\ dx$ $=$ $\frac{1}{2}\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n)+\frac{F(2)}{2}$
$=$ $\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n),$
from which we deduce that
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\int_{0}^{1}\big(\varphi_{f}(x+1)-f(x+1)\big)\ dx=\int_{0}^{1}\varphi_{f}(x)\ dx.$
Corollary 1.
Let $f\in\mathcal{E}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}$ and we have the shift formula
$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$ (12)
Proof.
Let $g:x\mapsto f(x+1)$, then it is easily verified, by theorem 3, that we have for $x>0$
$\varphi_{g}(x)=\varphi_{f}(x+1)-f(1).$
Thus we obtain
$\int_{0}^{1}\varphi_{g}(x)dx=\int_{0}^{1}\varphi_{f}(x+1)\ dx-f(1)=\int_{0}^{1}\varphi_{f}(x)\ \ dx+\int_{0}^{1}f(x+1)\ dx-f(1),$
and the conclusion follows from theorem 4. ∎
Example 4.
Let $f:z\mapsto 1/z$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\int_{0}^{1}(\Psi(x+1)+\gamma)\ \ dx=\int_{0}^{1}\frac{\Gamma^{\prime}}{\Gamma}(x+1)\ dx+\gamma,$
thus we verify again that
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\gamma,$ (13)
Observe that, by the shift formula (12), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}=\gamma-1+\log(2).$
Example 5.
Let $\log$ the principal determination of logarithm. Since, for any $x>0$, we have
$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1),$
and $\log(\Gamma(2))=\log(1)$, we deduce that $\varphi_{\log}(x)=\log(\Gamma(x+1))$. Now if we use the Raabe’s formula, that is :
$\int_{a}^{a+1}\log(\Gamma(x))\ dx=a\log(a)-a+\log(\sqrt{2\pi}),$
then we get
$\sum_{n\geq 1}^{\mathcal{R}}\log(n)=\log(\sqrt{2\pi})-1.$ (14)
Example 6.
Let $f=\Psi$ and let $\varphi:x\mapsto x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
$\varphi(x+1)-\varphi(x)=\Psi(x+1)=f(x+1).$
Thus, by the previous theorem, we have
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=\int_{0}^{1}x\Psi(x)\ dx+\frac{1}{2}=-\int_{0}^{1}\log(\Gamma(x))\ dx+\frac{1}{2},$
then by Raabe’s formula, we get
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$ (15)
Example 7.
We now evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for any $k$ integer $\geq 1$. Let the Bernoulli polynomials $B_{k}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
$B^{\prime}_{k}(x)=kB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)\ dx=0.$
The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. These polynomials verify (cf. [G-R])
$B_{k+1}\big((x+1)+1\big)-B_{k+1}(x+1)=(k+1)(x+1)^{k}.$
Let the function $f:x\mapsto x^{k}$ with $k$ integer $\geq 1$. Then by theorem 3, we obtain
$\varphi_{f}(x)=\frac{B_{k+1}(x+1)}{k+1}-\frac{B_{k+1}}{k+1}.$
Since
$\int_{0}^{1}\frac{B_{k+1}(x+1)}{k+1}\ dx=\int_{0}^{1}\frac{B_{k+1}(x)}{k+1}\ dx\,+\int_{0}^{1}x^{k}\ dx=\frac{1}{k+1},$
then theorem 4 gives
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}=\frac{1-B_{k+1}}{k+1}.$ (16)
Remark 5.
Relations (13) and (16) can be written in terms of the Zeta function :
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}$ $=$ $\lim_{s\to 1}\Big(\zeta(s)-\frac{1}{s-1}\Big),$
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ $=$ $\zeta(-k)-\frac{1}{-k-1}\text{ for every integer }k\geq 1.$
Note also that
$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1.$
These formulae are extensions of formula (8) and are special cases the following general result.
Theorem 5.
The function
$s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;$
is analytic on $\mathbb{C}$ and we have for every $s\in\mathbb{C}\setminus\{1\}$
$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}.$ (17)
Proof.
Let $s=\sigma+i\tau$ with $\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$, and $z=x+it$ with $x>\varepsilon>0$, $t\in\mathbb{R}$. We have
$\Big|\frac{1}{z^{s}}\Big|=\Big|\frac{1}{(x+it)^{s}}\Big|=\Big|e^{-s\log(x+it)}\Big|=e^{-{\sigma}\log\sqrt{x^{2}+t^{2}}+\tau\arctan(\frac{t}{x})}\leq e^{|\tau|\pi/2}(x^{2}+t^{2})^{-\sigma/2}$
Thus for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}$ and the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined. By (6), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt,$
$\frac{1}{2}+2\int_{0}^{+\infty}\frac{\sin(s\arctan(t))}{(1+t^{2})^{s/2}}\frac{1}{e^{2\pi t}-1}\ dt.$
This last integral depends analyticaly of $s\in\mathbb{C}$ so the function $s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is analytic on $\mathbb{C}$. Since the function $s\mapsto\zeta(s)-\frac{1}{s-1}$ is analytic on $\mathbb{C}\setminus\{1\}$ then, by analytic continuation of (8), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}\text{ for any }s\in\mathbb{C}\setminus\{1\}.$
3 An expansion of Zeta in the critical strip
We are now going to use expression (17) to obtain, in the strip $\{0<\Re(s)<1\}$, a polynomial expansion of Zeta
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s).$
3.1 The sequence of polynomials $(P_{m})$
For $m\geq 0$}, let $\psi_{m}$ the function defined in the half-plane $\{Re(z)>0\}$ by
$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m},$ (18)
and let $(P_{m}(s))$ is the sequence of polynomials given by the generating function
$\frac{2}{1-\xi}\Big({\frac{1-\xi}{1+\xi}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)\ \xi^{m}\ \text{ for }\ \ \ \xi\in D(0,1).$ (19)
Then, with $\xi=\frac{z-1}{z+1}$, we have for $s\in\mathbb{C}$ and $Re(z)>0$, the expansion
$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)\,P_{m}(s).$ (20)
To obtain an explicit expression of $P_{m}(s)$, let $\xi=\frac{x}{1+x}$, with $x\in]0,1[$, then by (19) we get
$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$
where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ Now the well-known equivalence (cf.[H1])
$\sum_{m\geq 0}b_{m}(\frac{x}{1+x})^{m+1}=\sum_{k\geq 0}a_{k}x^{k+1}\ \Leftrightarrow\ b_{m}=\sum_{k=0}^{m}\binom{m}{k}\ a_{k},$
gives for $P_{m}(s)$ the following expression
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{(s)_{k}}{k!}.$ (21)
It is easily verified that $P_{m}(s)$ is related, for $\Re(s)>0,$ with the Laguerre polynomial $L_{m}$ by
$P_{m}(s)=\frac{2}{\Gamma(s)}\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$ (22)
We have another integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
$\frac{(s)_{k}}{k!}=\frac{1}{\Gamma(s)\Gamma(1-s)}\int_{0}^{1}t^{s-1+k}(1-t)^{-s}\ dt,$
thus, by (21), we get for $0<\Re(s)<1$
$P_{m}(s)=2\frac{\sin(\pi s)}{\pi}\int_{0}^{1}t^{s-1}(1-t)^{-s}(1-2t)^{m}\ dt.$ (23)
This integral expression allows us to obtain the following lemma which will be useful later.
Lemma 1.
For $0<\Re(s)<1$, we have, for every $m\geq 1$
$|P_{m}(s)|\leq e^{\pi|Im(s)|}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big).$
Proof.
Using (23) we get
$|P_{m}(s)|$ $\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{0}^{1}t^{\Re(s)-1}(1-t)^{-\Re(s)}|1-2t|^{m}\ dt,$
$\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}\ du,$
$\leq$ $e^{\pi|Im(s)|}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}\ du.$
Let $a=\Re(s)$, this last integral is :
$I_{a}=\int_{0}^{1}(1-u)^{a-1}(1+u)^{-a}u^{m}\ du+\int_{-1}^{0}(1-u)^{a-1}(1+u)^{-a}(-u)^{m}\ du.$
Since $0<a<1$, we have
$I_{a}\leq\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{-1}^{0}(1+u)^{-a}(-u)^{m}\ du$ $=$ $\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{0}^{1}(1-u)^{-a}u^{m}\ du.$
Now if $0<\alpha<1$ we have
$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{m}\ du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-mt}\ dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-mt}\ dt=\Gamma(\alpha)\ m^{-\alpha},$
(the last inequality is justified by $e^{-t}\leq\frac{1-e^{-t}}{t}\leq{1}$, for every $t>0$). With $\alpha=a$ and $\alpha=1-a$, we obtain
$I_{a}\leq\frac{\Gamma(a)}{m^{a}}+\frac{\Gamma(1-a)}{m^{1-a}}.$
Remark 6.
The polynomials $P_{m}$ are related with Gauss hypergeometric function
$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!}.$
It is easily verified with (21) that, for $m\geq 1$, we have
$P_{m}(s)=2F(-m,s,1,2).$
By classical properties of hypergeometric function (cf. [G-R]), we get the relations
$P_{m}(1-s)=(-1)^{m}P_{m}(s),$ (24)
$sP_{m}(s+1)-(s-1)P_{m}(s-1)=(2m+1)P_{m}(s).$ (25)
3.2 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
Theorem 6.
For $0<\Re(s)<1,$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$ (26)
And we obtain the expansion
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}\,P_{m}(s).$ (27)
Proof.
We apply (20) with $z=n$ integer $\geq 1$, this gives
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\psi_{m}(n)P_{m}(s)=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$ (28)
We now prove that, for $0<\Re(s)<1$, these last two sums can be inverted, that is to say, the series
$\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$ (29)
is convergent for every $0<\Re(s)<1$, and its sum is equal to
$\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
By definition (6), we can write
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=c_{m}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt$
with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. Since $c_{0}P_{0}(s)=1/2$, then the series (29) is
$\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$
To prove the convergence of this series and invert signs $\sum$ and $\int$ we use the following lemma.
Lemma 2.
For $0<\Re(s)<1$, the series
$\sum_{m\geq 0}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|$ (30)
is convergent for every $t>0$, and its sum verify
$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}\ dt<+\infty.$
By this lemma the series (29) is convergent for every $0<\Re(s)<1$, and we get
$\sum_{m=0}^{+\infty}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\sum_{m=0}^{+\infty}\big(\frac{it}{2+it}\big)^{m}P_{m}(s)-\frac{1}{2-it}\sum_{m=0}^{+\infty}\big(\frac{-it}{2-it}\big)^{m}P_{m}(s)\Big)\frac{1}{e^{2\pi t}-1}\ dt.$
Using (19) with $\xi=\frac{\pm it}{2\pm it}$, this last expression is equal to
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
The relation (27) follows immediately from (17).
Proof of lemma 2.
We have
$\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big|\leq 2\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}\Big|=\frac{2}{\sqrt{t^{2}+4}}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}$
Using lemma 1, we see that the terms of the series (30) are, for $m\geq 1$, dominated by the terms
$\frac{4}{\sqrt{t^{2}+4}}e^{\pi|Im(s)|}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
Consequently, the series (30) is convergent, and we have
$\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \leq M(t)$
with
$M(t)=\frac{4t}{4+t^{2}}+e^{\pi|Im(s)|}\frac{4}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
It remains to prove that
$\int_{0}^{+\infty}M(t)\frac{1}{e^{2\pi t}-1}\ dt<+\infty.$
This follows from the convergence of the integral
$\int_{0}^{+\infty}\frac{4t}{4+t^{2}}\frac{1}{e^{2\pi t}-1}\ dt,$
and the fact that, for $0<\alpha<1$, we have
$\int_{0}^{+\infty}\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\frac{1}{e^{2\pi t}-1}\ dt<+\infty,$
as a consequence of the following inequalities
$\frac{1}{\sqrt{t^{2}+4}}\sum_{m=2}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\leq\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\leq\frac{t}{{t^{2}+4}}\Big(\frac{1}{1-\frac{t}{\sqrt{t^{2}+4}}}\Big)\leq\frac{t^{2}}{4}.$
3.3 Evaluation of the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
Lemma 3.
For any integer $m\geq 0$
$z_{m}:=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=x_{m}+y_{m}$
with $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\Big(\zeta(k+1)-\frac{1}{k}\Big)-(-1)^{m},$ (31)
and
$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx=\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$ (32)
Proof.
We apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, then we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
For any $m$ integer $\geq 0$, let
$x_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}\ \text{ and }\ y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
We have $x_{0}=\gamma-1,\ \ y_{0}=\log(2).$ And by the binomial expansion
$\Big(1-\frac{2}{x}\Big)^{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\frac{1}{x^{k}}$
we get for $m\geq 1$
$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}-(-1)^{m}.$
and
$y_{m}=\int_{1}^{2}\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}\ \ dx=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}\ dx.$
The conclusion follows from (17) which gives for $k\geq 1$
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}=\zeta(k+1)-\frac{1}{k}.$
3.4 An expansion of Zeta
Theorem 7.
For $0<\Re(s)<1$, we have
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}\,P_{m}(s),$ (33)
where
$z_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\frac{1}{k}.$
Proof.
This follows from the expansion (27) of theorem 6. Since by lemma 3, we have
$z_{m}=x_{m}+y_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k},$
and we can use the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1}{k}.$
The previous theorem will enable us to get an expansion of Zeta in the strip $\{0<\Re(s)<1\}$. Indeed, the polar part 1/(s1)1/(s-1) of (33) has the simple expansion given by the following lemma.
Lemma 4.
For $0<\Re(s)<1$ we have
$\frac{1}{s-1}=-\sum_{m=0}^{+\infty}y_{m}\,P_{m}(s).$ (34)
Proof.
For $0<\Re(s)<1$, we write
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\ dx.$ (35)
This interchanging of signs $\sum$ and $\int$ is justified by
$\sum_{m=0}^{+\infty}\int_{0}^{1}|\frac{x-1}{x+1}|^{m}\ |P_{m}(s)|\frac{1}{x+1}\ dx\leq\sum_{m=0}^{+\infty}\int_{0}^{1}\frac{1}{(x+1)^{m+1}}\ dx\ |P_{m}(s)|\leq\sum_{m=0}^{+\infty}\frac{1}{m}|P_{m}(s)|,$
and we immediately check, with lemma 1, that the latter series is convergent.
The relation (34) is then a consequence of (20) and (35).
Theorem 8.
For $0<\Re(s)<1$, we have
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}\,P_{m}(s)$ (36)
with $x_{0}=\gamma-1,$ and for $m\geq 1$
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$ (37)
and $r_{m}$ are the rational numbers
$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
Proof.
This is an immediate consequence of the preceding lemma and theorem 7 and the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-2)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
For the first values of $m$ we get
$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3$
$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
3.5 Orthogonality relations
The polynomial $P_{m}$ is related to the Mellin transform of the function
$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x),$ (38)
since we verify easily with (22) that, for $\Re(s)>0$, we have
$\mathcal{M}\varphi_{m}(s)=\int_{0}^{+\infty}\varphi_{m}(x)x^{s-1}\ \ dx=\frac{1}{\sqrt{2}}{\Gamma(s)}P_{m}(s).$ (39)
It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),\ dx)$.
Theorem 9.
The sequence $(Q_{m})_{m\geq 0}$ of polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$ (40)
is an orthonormal system in $\big(L^{2}(\mathbb{R}),\frac{dt}{\cosh(\pi t)}\big)$. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
Proof.
By the Mellin-Parseval formula (cf. [T]), we have
$\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(\varphi_{n})(\frac{1}{2}+it)\ dt=\int_{0}^{+\infty}\overline{\varphi_{m}(x)}\varphi_{n}(x)\ dx=\delta_{(m,n)},$
with $\delta$ the usual Kronecker symbol. That is
$\frac{1}{4\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)}\overline{P_{m}(\frac{1}{2}+it)}\ \Gamma(\frac{1}{2}+it)P_{n}(\frac{1}{2}+it)\ dt=\delta_{(m,n)}.$
Since
$\Gamma(\frac{1}{2}+it)\overline{\Gamma(\frac{1}{2}+it)}=\frac{\pi}{\cosh\pi t},$
then
$\frac{1}{4}\int_{-\infty}^{+\infty}\overline{P_{m}(\frac{1}{2}+it)}\ P_{n}(\frac{1}{2}+it)\frac{dt}{\cosh\pi t}=\delta_{(m,n)}.$
By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. Thus the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. ∎
Remark 7.
By (24) we have, for every $m\geq 0$, the relations
$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
thus we have for $t\in\mathbb{R}$
$\Re\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m}Q_{2m}(t)$
$\Im\zeta(\frac{1}{2}+it)$ $=$ $-2i\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$
By (25) we have,
$TQ_{m}=(2m+1)Q_{m},$
where $T$ is the operator on $\mathbb{R}[t]$ defined by
$Tf(t)=\big(\frac{1}{2}+it\big)f(t-i)+\big(\frac{1}{2}-it\big)f(t+i).$
3.6 An orthonormal expansion
Let $X(0)=\gamma-1$ and for $k\geq 1$ let
$X(k)=\zeta(k+1)-1-\frac{1}{k}.$
Then formula (31) can be written as a binomial transform
$x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$ (41)
Let the function $f_{0}$ defined for $t>0$ by
$f_{0}(t)=\frac{1}{e^{t}-1}-\frac{1}{t},$
it is well-known that $\mathcal{M}f_{0}(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$. We easily verify that for $k\geq 0$ we have
$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)\ dt.$
Then by definition of Laguerre polynomials, we obtain
$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$
This means that we have, in $L^{2}(]0,+\infty[)$, the orthonormal expansion
$f_{0}=\sqrt{2}\,\sum_{m=0}^{+\infty}x_{m}\,\varphi_{m}.$ (42)
Theorem 10.
For any integer $m\geq 0$, we have
$x_{m}=\frac{1}{2}\int_{-\infty}^{+\infty}\overline{Q_{m}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$ (43)
Proof.
By (42) and Mellin-Parseval formula (cf. [T]), we have
$x_{m}=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)\ dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt,$
thus we get
$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \Gamma((\frac{1}{2}+it))\zeta(\frac{1}{2}+it)\ dt$
$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Corollary 2.
We have
$\gamma=1+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)},$
and for $m\geq 1$
$\zeta(m+1)=1+\frac{1}{m}+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{(\frac{1}{2}-it)_{m}}{m!}\frac{dt}{\cosh(\pi t)}.$
Proof.
With (41) and inversion of the binomial transform, we get
$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x_{k}=2^{m}X(m),$
and from (21)
$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{\frac{1}{2}P_{k}(\frac{1}{2}+it)}=2^{m}\ \frac{(\frac{1}{2}-it)_{m}}{m!}.$
Thus by (43) we obtain the integral expression
$X(m)$ $=$ $\frac{1}{2^{m+1}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\int_{-\infty}^{+\infty}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$=$ $\frac{1}{2^{m+1}}\int_{-\infty}^{+\infty}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\frac{(\frac{1}{2}-it)_{m}}{m!}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Example 8.
For $m=1,2,$ we have
$\zeta(2)$ $=$ $2+\frac{1}{2}\int_{-\infty}^{+\infty}(\frac{1}{2}-it)\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
$\zeta(3)$ $=$ $\frac{3}{2}-\frac{1}{4}\int_{-\infty}^{+\infty}(t^{2}+2it-\frac{3}{4})\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Remark 8.
The previous theorem means that the pointwise convergent expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}\,Q_{m}(t)$
is also an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Then we have
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=4\sum_{m=0}^{+\infty}x_{m}^{2}.$ (44)
We now give a formal evaluation of this last sum using Ramanujan’s summation. Let $\mathcal{L}$ the Laplace transform
$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx.$
We have (cf. [G-R]) for $Re(z)>0$, the Laplace transforms
$\mathcal{L}\varphi_{m}(z)$ $=$ ${\sqrt{2}}\ \psi_{m}(z),$
$\mathcal{L}f_{0}(z)$ $=$ $\log(z)-\Psi(z)-\frac{1}{z}.$
By Laplace transform of expansion (42) we get
$2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(z)=\log(z)-\Psi(z)-\frac{1}{z}.$ (45)
Now, if we apply (45) with $z=1,2,3,\dots$ and take Ramanujan’s summation, we get
$2\sum_{m=0}^{+\infty}x_{m}\sum_{n\geq 1}^{\mathcal{R}}\psi_{m}(n)=2\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}x_{m}\psi_{m}(n)=\sum_{n\geq 1}^{\mathcal{R}}\log(n)-\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)-\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}.$
Then, with the evaluations of these sums given in lemma 3 and in (13), (14), (15), we get
$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{m})=\log(2\pi)-\gamma-\frac{3}{2}.$
Since by (32) $y_{m}=\int_{0}^{1}\psi_{m}(x)\ dx$ then using (45) we have
$2\sum_{m=0}^{+\infty}x_{m}y_{m}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)\ dx=\int_{0}^{1}(\log(x)-\Psi(1+x))\ dx=-1.$
Finally we get
$\sum_{m=0}^{+\infty}x_{m}^{2}=\frac{1}{2}\log(2\pi)-\frac{1}{2}\gamma-\frac{1}{4}.$
With (44) this gives
$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$ (46)